A diffraction grating with 600 lines/mm is illuminated with light of wavelength 510 nm. A very wide viewing screen is 4.2 m behind the grating.

Part A

What is the distance between the two m = 1 bright fringes?

Express your answer with the appropriate units.

Δy =

SubmitMy AnswersGive Up

Part B

How many bright fringes can be seen on the screen?

N =

SubmitMy AnswersGive Up
Answer #1

(A) d = 1 x 10^-3 m / 600 = 1.67 x 10^-6


y = m lambda D / d

y = (1 x 510 x 10^-9 x 4.2) / (1.67 x 10^-6)

y = 1.2852 m


d = 2 y = 2.57 m

(B) sin(theta) = m lambda / d


sin90 = m (510 x 10^-9) / (1.67 x 10^-6)

m = 3.26


total bright fringe = 2(3) + 1 = 7